In a shipment of 11 CD players, two are defective. Find the probability that a sample of 3 CD players will contain no more than one defective one.
Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

In a shipment of 11 CD players, two are defective. Find the probability that a sample of 3 CD players will contain no more than one defective one.
------------------
P(defective) = 2/11
P(not defective = 9/11
----------------
Binomial Problem with n = 3.
--
P(0<= x <=1) = P(x=0) + P(x=1)
---
= (9/11)^3 + 3(2/11)(9/11)^2
---
= 0.5477 + 0.3651
---
= 0.9128
====================
Cheers,
Stan H.
=============================
Cheers,
Stan H.
============

Answer by edjones(8007)

(9C1*2C2)/11C3
=9/165 probability that sample contains 2 defective players.
1 - 9/165 = 156/165 the probability that a sample of 3 CD players will contain no more than one defective one.
.
Ed